# Socioeconomic Drivers of Physical Morbidity Across U.S. Counties: A Spatial Causal Inference Approach

### 28 Mar 2026 v1

@@ 3.2 Spatial Basis Construction
Eigenvectors of $\widetilde{\mathbb{W}}$ associated with the largest eigenvalues maximize Moran’s $I$ statistic, and also isolate patterns of positive spatial autocorrelation [griffith2021interpreting]. We use the first $k$ eigenvectors form $\mathbf{Z}_{\text{MEM}}\in\mathbb{R}^{n\times k}$ as the basis function design matrix for our problem.
@@
(Intrinsic) Conditional Autoregressive (ICAR) models specify spatial dependence via the graph Laplacian precision matrix $\mathbb{Q}=\mathbb{D}-\rho\mathbb{W}$, where $\mathbb{D}$ is the diagonal degree matrix, and $\rho$ is the autocorrelation parameter – set as $1$ for the intrinsic problems. Since this is the precision matrix, eigenvectors corresponding to the smallest non-zero eigenvalues of $\mathbb{Q}$ capture smooth, large-scale spatial structure. We retain the first $k$ eigenvectors (excluding the constant null vector), forming $\mathbf{Z}_{\text{ICAR}}\in\mathbb{R}^{n\times k}$.
Both bases provide orthogonal decompositions of spatial variation, as they are derived from eigenvectors of spatial covariance (MEM) or precision (ICAR) matrices. We set $k=400$ based on residual spatial autocorrelation diagnostics (Moran’s $I$ test, refer to Table 1). To avoid overfitting, we penalize the regression coefficients on these basis functions in the causal estimation procedure described below.
@@ 4 Results
Table 2 reports doubly robust GPS estimates for six socioeconomic and environmental determinants, all of which pass the residual autocorrelation check (Moran’s $I$, $p>0.05$). Food insecurity and unemployment rate emerge as the dominant drivers of physical morbidity. A one percentage point increase in food insecurity corresponds to 0.086 additional unhealthy days per month, so a five point increase implies roughly 5.2 additional days per year. The unemployment rate effect is also similar in magnitude: a five point increase corresponds to approximately 3.1 additional unhealthy days per year. Pay gap carries a modest but statistically significant positive effect. Homeownership is borderline significant in the negative direction, plausibly reflecting its role as a broad proxy for household economic stability. Rent burden and PM 2.5 show no significant effects after spatial adjustment – a finding that may reflect weak marginal relationships at the county scale, attenuation from geographic aggregation, or both.

### 31 Mar 2026 v2

@@ 3.2 Spatial Basis Construction
Eigenvectors of $\widetilde{\mathbb{W}}$ associated with the largest eigenvalues maximize Moran’s II statistic, and also isolate patterns of positive spatial autocorrelation [griffith2021interpreting]. We use the first $k$ eigenvectors form $\mathbf{Z}_{\text{MEM}}\in\mathbb{R}^{n\times k}$ as the basis function design matrix for our problem.
@@
(Intrinsic) Conditional Autoregressive (ICAR) models specify spatial dependence via the graph Laplacian precision matrix $\mathbb{Q}=\mathbb{D}-\rho\mathbb{W}$, where $\mathbb{D}$ is the diagonal degree matrix, and $\rho$ is the autocorrelation parameter – set as 11 for the intrinsic problems. Since this is the precision matrix, eigenvectors corresponding to the smallest non-zero eigenvalues of $\mathbb{Q}$ capture smooth, large-scale spatial structure. We retain the first $k$ eigenvectors (excluding the constant null vector), forming $\mathbf{Z}_{\text{ICAR}}\in\mathbb{R}^{n\times k}$.
Both bases provide orthogonal decompositions of spatial variation, as they are derived from eigenvectors of spatial covariance (MEM) or precision (ICAR) matrices. We set $k=400$ based on residual spatial autocorrelation diagnostics (Moran’s II test, refer to Table 1). To avoid overfitting, we penalize the regression coefficients on these basis functions in the causal estimation procedure described below.
@@ 4 Results
Table 2 reports doubly robust GPS estimates for six socioeconomic and environmental determinants, all of which pass the residual autocorrelation check (Moran’s II, $p>0.05$). Food insecurity and unemployment rate emerge as the dominant drivers of physical morbidity. A one percentage point increase in food insecurity corresponds to 0.086 additional unhealthy days per month, so a five point increase implies roughly 5.2 additional days per year. The unemployment rate effect is also similar in magnitude: a five point increase corresponds to approximately 3.1 additional unhealthy days per year. Pay gap carries a modest but statistically significant positive effect. Homeownership is borderline significant in the negative direction, plausibly reflecting its role as a broad proxy for household economic stability. Rent burden and PM 2.5 show no significant effects after spatial adjustment – a finding that may reflect weak marginal relationships at the county scale, attenuation from geographic aggregation, or both.